OpenStudy (anonymous):

What is the solution set for the equation |2x-3|=6? please explain how you get your answer

OpenStudy (freckles):

|2x-3|=6 means you have two equations to solve: 2x-3=6 or 2x-3=-6

OpenStudy (freckles):

in general if a is positive and you have |f(x)|=a then we have f(x)=a or f(x)=-a.

OpenStudy (anonymous):

ok

OpenStudy (freckles):

anyways all you have to do now is solve the following two linear equations: 2x-3=6 , 2x-3=-6

OpenStudy (freckles):

can you do that?

OpenStudy (freckles):

solving 2x-3=6 is only 2 steps solving 2x-3=-6 is also only 2 steps

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so for 2x-3=6 x=4.9 and for 2x-3=-6 x=-1.5?

OpenStudy (freckles):

9/2 should be 4.5

OpenStudy (freckles):

but yes the solution you have to the last equation you wrote is right

OpenStudy (anonymous):

ok

OpenStudy (freckles):

the solution set is {9/2,-3/2} or if you want to write in decimal form {4.5,-1.5}

OpenStudy (anonymous):

oh ok thank you!!!